OpenStudy (anonymous):

Evaluate the following indefinite integral Just to double check my answer @hartnn

OpenStudy (anonymous):

\[\int\limits_{}^{}(\sqrt{x}-\frac{ 1 }{ \sqrt{x} })dx\]

OpenStudy (anonymous):

my final answer is\[\frac{ 2 }{ 3 }x^{3/2}-2\sqrt{x}+c\]

hartnn (hartnn):

yup, thats correct! good work MARC :)

OpenStudy (anonymous):

Thank you @hartnn ^_^